# On Modular Invariants of Truncated Polynomial Rings in low ranks

### v1

@@ 8. A filtration of submodules
For $n=3$, we have shown that this increasing filtration of subspaces is exhaustive. The calculation in previous sections shows that our basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{G}(3)$ is natural in the sense that the resulting filtration is a filtration of ${\mathcal{D}}_{3}$-submodules. Moreover, the action of the Dickson invariants is very similar to the pattern discussed in [10] about the structure of the cofixed space $S_{G}$ as a $S^{G}$-module. Indeed, we can easily check that $Q_{3,0}$ annihilates $\mathcal{F}_{3,2}$; $Q_{3,0}$ and $Q_{3,1}$ annihilate $\mathcal{F}_{3,1}$; and $Q_{3,0},Q_{3,1},Q_{3,2}$ annihilate $\mathcal{F}_{3,0}$. A similar pattern also occurs in the case $n=2$.
The aim of this final section is to demonstrate Theorem 1.11. We will show the filtration $\mathcal{F}_{n,*}$ also behaves well with respect to the action of the mod $q$ Steenrod algebra $\mathcal{A}$, at least in the case $n\leq 3$.
@@
For background on the mod $q$ Steenrod algebra and its role in modular invariant theory, we refer to Larry Smith’s article [16], which treats the Steenrod algebra completely algebraically over a general finite field.
@@
Recall that for each $i\geq 0$, the $i$th Steenrod reduced power operation $\mathcal{P}^{i}$ is a $\mathbb{F}_{q}$-linear map satisfying the following conditions:
@@
Moreover, $\mathcal{P}^{0}$ acts as the identity. Over the polynomial ring $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, in which $x_{i}$ are given degree $1$, the action of the Steenrod algebra is completely determined by the unstable condition and the Cartan formulae. More concretely, we have
@@
for any linear polynomial $v$. Here the binomial coefficient is taken modulo $p$. Thus, if $j$ is a $q$-power, then $\mathcal{P}^{i}(v^{j})\neq 0$ iff $i=0$ (in which case $\mathcal{P}^{0}$ is the identity operator) or $i=j$ (in which case $\mathcal{P}^{i}(v^{i})=v^{qi}$ is the Frobenius operator).
The Steenrod algebra action and the $G$ action on $S$ commutes. Furthermore, it is clear that $I_{m}$ is stable under the action of the Steenrod operations, so is ${\mathcal{Q}}_{m}=S/I_{m}$. It follows that a Steenrod reduced power yields a map from ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ to itself. In other words, ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ is an $\mathcal{A}$-module.
The proof of Theorem 1.11 uses several lemmas. We begin with a general formula describing how the Steenrod operation interacts with the delta operator.
@@
If $f$ is a polynomial which is $\operatorname{GL}_{1}$-invariant in the first variable and $1\leq k\leq\deg(\delta_{2}(f))=q^{m}-q+\deg(f)$, then
@@
Moreover, the last two summands on the right-hand side vanish in ${\mathcal{Q}}_{m}$.
@@
The condition imposed on $f$ assures that $\delta_{2}(f)$ as well as the other terms in this lemma are polynomial. The equation above is then obtained by applying the Cartan formula for the product $L_{2}\delta_{2}(f)$ and then dividing both sides by $L_{2}$. Note that $\mathcal{P}^{q}L_{2}=Q_{2,1}L_{2}$, $\mathcal{P}^{q+1}L_{2}=Q_{2,0}L_{2}$ and $\mathcal{P}^{k}L_{2}=0$ for all $k\geq 1$. ∎
@@
For each $0\leq k\leq 2$, $\mathcal{F}_{2,k}$ is a submodule of ${\mathcal{Q}}_{m}^{G}(2)$ over the Steenrod algebra.
@@
We need to compute the action of the Steenrod operation on a typical element $\delta_{2}(f)$ where $f=Q_{1,0}^{a}$, $a<[m]_{q}$. In the rank $2$ case, the last two summands on the right-hand side of the formula given in Lemma 8.1 vanish in ${\mathcal{Q}}_{m}$. Indeed, one check by direct inspection that the degree of $\delta_{2;m+1}(Q_{1,0}\mathcal{P}^{k-1}f)$ is strictly greater than $2(q^{m}-1)$ for all $q$ and $m$. For $\delta_{2;m+1}(\mathcal{P}^{k-q^{m}}f)$, its degree is not greater than $2(q^{m}-1)$ only when $q=2$, $k\geq q^{m}$ and $\deg f=0$, but in this case $k>q^{m}-q+\deg(f)$.
Now we can proceed by induction on $k$ using the above formula and the calculation in Proposition 2.5. ∎
Corollary 8.3.
$\mathcal{F}_{3,1}$ is a submodule of ${\mathcal{Q}}_{m}^{G}(3)$ over the Steenrod algebra.
@@
We need to show that for each $k\geq 1$ and $a<[m]_{q}$, $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ is in the span of $\delta_{1}^{3}(1)$ and $\delta_{2}^{2}(\Delta^{m}_{1})$.
Apply the formula provided in Lemma 8.1 for $f=\delta_{2}Q_{1,0}^{a}$. Again, the last two summands are trivial in ${\mathcal{Q}}_{m}$ because of the instability condition and degree calculation. We then obtain a recursive equation in ${\mathcal{Q}}_{m}$:
$\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}Q_{1,0}^{a})+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}Q_{1,0}^{a})=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}\delta_{2}Q_{1,0}^{a})+\delta_{2}(\mathcal{P}^{k}\delta_{2}Q_{1,0}^{a}).$
We now proceed by double induction on $k$ that $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ belongs to $\mathcal{F}_{3,1}$. We already know from Corolarry 8.2 that $\mathcal{P}^{k-1}(\delta_{2}Q_{1,0}^{a})$ and $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{a})$ belongs to the span of $\delta_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}(1)$. By Lemma 7.2, the two summands on the right-hand side in the equation above belong to the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$.
Also by induction, we see that $Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}Q_{1,0}^{a})=0$ in ${\mathcal{Q}}_{m}$ and $Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}Q_{1,0}^{a})$ is also in the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$. We conclude that $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{a})$ is also in the span of $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{2}({\mathcal{Q}}_{m}(1))$, but since it is a $\operatorname{GL}_{3}$-invariant, it is spanned by $\delta^{2}_{2}(\Delta^{m}_{1})$ and $\delta_{1}^{3}(1)$. ∎
@@
If $f\in{\mathcal{D}}_{2}$ and $k\geq 1$, then
$\mathcal{P}^{k}(\delta_{3}f)+\sum_{i<k}\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)=\frac{\mathcal{P}^{k}\begin{vmatrix}x_{1}&x_{2}&x_{3}\\ x_{1}^{q}&x_{2}^{q}&x_{3}^{q}\\ x_{1}^{q^{m}}f(x_{2},x_{3})&x_{2}^{q^{m}}f(x_{1},x_{3})&x_{3}^{q^{m}}f(x_{1},x_{2})\end{vmatrix}}{L_{3}}.$
Moreover, the right-hand side equals $\delta_{3;m}(F)$ or $\delta_{3;m+1}(F)$ where $F$ is a certain linear combination of polynomials of the form $g\mathcal{P}^{i}f$ where $g\in{\mathcal{D}}_{2}$ and $0\leq i\leq k$.
@@
The above equality is a simple application of the Cartan formula for $\mathcal{P}^{k}(L_{2}\delta_{2;m}(f))$ (or $\mathcal{P}^{k}(L_{3}\delta_{3}f)$). For the right-hand side, note that because of the instability condition, we have $\mathcal{P}^{i}(x_{1})$ is either $x_{1}$ or $x_{1}^{q}$, depending on whether $i=0$ or $i=1$, and there is a similar formula for other entries in the determinant.
In the first case, $\delta_{2;m+1}(g\mathcal{P}^{i}f)$ is trivial in ${\mathcal{Q}}_{m}$ unless $q=2$ and $g\mathcal{P}^{i}f=1$. In this case, $\delta_{2;m+1}(1)=x_{1}^{2^{m}-1}x_{2}^{2^{m}-1}$ because of degree reason. We also have $g=f=1$, and $k=2^{m}$, but then $\delta_{2}f$ has degree $2^{m}-2<2^{m}$.
Finally, observe that any determinant of the form
$\begin{vmatrix}x_{1}^{q^{a}}&x_{2}^{q^{a}}&x_{3}^{q^{a}}\\ x_{1}^{q^{b}}&x_{2}^{q^{b}}&x_{3}^{q^{b}}\\ x_{1}^{q^{c}}\mathcal{P}^{d}f(x_{2},x_{3})&x_{2}^{q^{c}}\mathcal{P}^{d}f(x_{1},x_{3})&x_{3}^{q^{c}}\mathcal{P}^{d}f(x_{1},x_{2})\end{vmatrix}$
can be written as
$L_{3}\delta_{3;c}\bigg{(}\frac{[a,b]}{[0,1]}\mathcal{P}^{d}f\bigg{)}.$
where $\frac{[a,b]}{[0,1]}\in\mathcal{D}_{2}$. Note that $c$ is either $m$ or $m+1$. ∎
It is also clear that $\mathcal{P}^{i}L_{3}$ is again a matrix whose entries are $q$-powers, so it is always a multiple of $L_{3}$ and the quotient $\frac{\mathcal{P}^{i}L_{3}}{L_{3}}$ is a Dickson polynomial in ${\mathcal{D}}_{3}$. The next lemma discusses how to deal with the case when $c=m+1$.
@@
The degree of $\delta_{3;m+1}(h)$ equals $q^{m+1}-q^{2}+\deg(h)$. If $q>3$, then this is strictly greater than $3(q^{m}-1)$, and we immediately have $\delta_{3;m+1}(h)\equiv 0$ in this case.
From the fundamental equation, we know that $X^{q^{3}}-Q_{3,2}X^{q^{2}}+Q_{3,1}X^{q}-Q_{3,0}X=0$ if $X$ is in the span of $x_{1},x_{2},x_{3}$. It follows that
$x_{i}^{q^{m+1}}-Q_{3,2}^{q^{m-2}}x_{i}^{q^{m}}+Q_{3,1}^{q^{m-2}}x_{i}^{q^{m-1}}-Q_{3,0}^{q^{m-2}}x_{i}^{q^{m-2}}=0,$
and so
@@
If $q=3$, then $q^{m+1}-q^{2}+\deg(h)\leq 3(q^{m}-1)$ implies $\deg(h)\leq q^{2}-3=6$. Since $h\in{\mathcal{D}}_{2}$, it follows that up to a scalar, $h=1$ or $h=Q_{2,1}$.
In the case where $h=Q_{2,1}$, $\delta_{3;m+1}(Q_{2,1})$ has degree $33^{m+1}-$ 3 so that it is either zero or the top element $x_{1}^{3^{m}-1}x_{2}^{3^{m}-1}x_{3}^{3^{m}-1}$. We claim that $\delta_{3;m+1}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m}$. Indeed, when $q=3$, then by direct computation, we have $Q_{3,0}=Q_{3,1}=0$ in ${\mathcal{Q}}_{2}$ so that $Q_{3,1}^{q^{m-2}}=Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}$. It follows that in ${\mathcal{Q}}_{m}$, we have
$\delta_{3;m+1}(Q_{2,1})=Q_{3,2}^{q^{m-2}}\delta_{3;m}(Q_{2,1}).$
On the other hand, $Q_{3,2}=x_{1}^{6}x_{2}^{6}x_{3}^{6}$ in ${\mathcal{Q}}_{2}$ so that $Q_{3,2}^{q^{m-2}}=x_{1}^{2.3^{m-1}}x_{2}^{2.3^{m-1}}x_{3}^{2.3^{m-1}}$ in ${\mathcal{Q}}_{m}$. So if we can show that $\delta_{3;m}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m-1}$, then it follows immediately that $\delta_{3;m+1}(Q_{2,1})=0$ in ${\mathcal{Q}}_{m}$. The claim can now be easily verified by checking that $\delta_{3;3}(Q_{2,1})=Q_{3,1}=0$ in ${\mathcal{Q}}_{2}$.
Suppose now $h=1$. In this case, we have
$\delta_{3;m+1}(1)=Q_{3,2}^{3^{m-2}}\delta_{3;m}(1)=\delta_{3;m}(Q_{2,1}^{3^{m-1}}),$
as required. Here the second equality follows from 2.5.
Now consider the case $q=2$. We still have $Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}(3)$, but now
$Q_{3,1}^{2^{m-2}}=x_{1}^{2^{m-1}}x_{1}^{2^{m-1}}x_{1}^{2^{m-1}}.$
By direct inspection, we have $\delta_{3;m+1}(h)=Q_{3,2}^{2^{m-2}}\delta_{3;m}(h)-Q_{3,1}^{2^{m-2}}\delta_{3;m-1}(h)$ equals in ${\mathcal{Q}}_{m}$ to
@@
It is straightforward to verify that
@@
so the expression above becomes
$\delta_{3;m}(Q_{2,1}h\delta_{2;m}(1))-\delta_{3;m}(h\delta_{2;m}(Q_{1,0}^{2})).$
Now we can use the argument as in the proof of Lemma 7.3 to conclude. ∎
@@
With a little more work, using 7.2, it can be showed that when $q=2$, then for $h\in{\mathcal{D}}_{2}$, we have an explicit formula:
$\delta_{3;m+1}(h)=\begin{cases}0&\text{if $h$ is a multiple of $Q_{2,0}$,}\\ \delta_{2;m}^{2}(Q_{1,0}^{2s})&\text{if $h=Q_{2,1}^{s}$}.\end{cases}$
@@
Now we can finish the proof of Theorem 1.11. It remains to show that $\mathcal{F}_{3,1}$ is an $\mathcal{A}$-submodule of ${\mathcal{Q}}_{m}^{G}(3)$. From Lemma 8.5 we see that if $f\in{\mathcal{D}}_{2}$, then
$\mathcal{P}^{k}(\delta_{3}f)+\sum_{i<k}\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)$
belongs to $\mathcal{F}_{3,1}$. By induction, so is $\mathcal{P}^{i}(\delta_{3}f)$ for each $i<k$ and from Proposition 2.5, $\frac{\mathcal{P}^{k-i}L_{3}}{L_{3}}\mathcal{P}^{i}(\delta_{3}f)$ also belongs to $\mathcal{F}_{3,1}$.

### v2

@@ 8. A filtration of submodules over the Steenrod algebra
The calculations in the previous sections demonstrate that our basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}(3)^{G}$ is natural in the sense that the resulting filtration is a filtration by ${\mathcal{D}}_{3}$-submodules. Moreover, the action of the Dickson algebra follows a pattern similar to the one discussed in [9], where the structure of the cofixed space $S_{G}$ is described as an $S^{G}$-module. A similar pattern also emerges in the case $n=2$.
@@
We now turn our attention to the Steenrod algebra. For background on the mod $q$ Steenrod algebra $\mathcal{A}$ and its role in modular invariant theory, we refer to Larry Smith’s article [13], which provides a purely algebraic treatment of the Steenrod algebra over a general finite field.
@@
Recall that for each $i\geq 0$, the $i$th Steenrod reduced power operation $\mathcal{P}^{i}$ is an $\mathbb{F}_{q}$-linear map satisfying the following conditions:
@@
Moreover, $\mathcal{P}^{0}$ acts as the identity. Over the polynomial ring $S=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$, in which each $x_{i}$ is given degree $1$, the action of the Steenrod algebra is completely determined by the unstable condition and the Cartan formulae. More concretely, we have
@@
for any linear polynomial $v$. Here the binomial coefficient is taken modulo $p$. Thus, if $j$ is a $q$-power, then $\mathcal{P}^{i}(v^{j})\neq 0$ iff $i=0$ (in which case $\mathcal{P}^{0}$ is the identity operator) or $i=j$ (in which case $\mathcal{P}^{i}(v^{i})=v^{qi}$ is the Frobenius operator). It is convenient to make the convention that $\mathcal{P}^{k}=0$ if $k<0$.
The Steenrod algebra action and the group $G$ action on $S$ commute. Moreover, it is evident that the Frobenius ideal $I_{m}$ is stable under the action of the Steenrod operations. Consequently, the quotient ${\mathcal{Q}}_{m}=S/I_{m}$ inherits the structure of an $\mathcal{A}$-module. It follows that each Steenrod reduced power $\mathcal{P}^{i}$ induces a well-defined endomorphism of ${\mathcal{Q}}_{m}^{G}$. In other words, the subspace of invariants ${\mathcal{Q}}_{m}^{G}$ is closed under the action of the Steenrod algebra, and therefore forms an $\mathcal{A}$-module.
The proof of Theorem 1.12 uses several lemmas. We begin with a general formula describing how the Steenrod operations interact with the delta operator.
@@
If $f$ is a polynomial which is $G_{1}$-invariant in the first variable, then for any $k\geq 0$, the following holds in $S$:
@@
The condition imposed on $f$ assures that $\delta_{2}(f)$ as well as the other terms in this lemma are polynomial. The equation above is then obtained by applying the Cartan formula for the product $L_{2}\delta_{2}(f)$ and then dividing both sides by $L_{2}$. Note that $\mathcal{P}^{q}L_{2}=Q_{2,1}L_{2}$, $\mathcal{P}^{q+1}L_{2}=Q_{2,0}L_{2}$ and $\mathcal{P}^{k}L_{2}=0$ for all other $k\geq 1$. ∎
@@
$\mathcal{F}_{2,1}$ is a submodule of ${\mathcal{Q}}_{m}(2)^{G}$ over the Steenrod algebra.
@@
We must show that for each $0\leq s<[m]_{q}$, $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{s})\in\mathcal{F}_{2,1}$ for all $k>0$. If $k\geq q^{m}$, the degree of $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{s})$ is at least
$q^{m}(q-1)+q^{m}-q+s(q-1)\geq 2(q^{m}-1).$
So we can assume that $k<q^{m}$. In this case, when we apply the formula of Lemma 8.3 for $f=Q_{1,0}^{a}$, the last two summands on the right-hand side vanish. Thus we have an equality in $S$:
(8.1) $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{s})+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}Q_{1,0}^{s})+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}Q_{1,0}^{s})=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}Q_{1,0}^{s})+\delta_{2}(\mathcal{P}^{k}Q_{1,0}^{s}).$
Both summands on the right-hand side are evidently of the form $\delta_{2}(Q_{1,0}^{s})$. Now we can proceed by induction on $k$ using the above formula and the calculation in Proposition 2.7 when working in ${\mathcal{Q}}_{m}(2)$. ∎
Note that $\mathcal{F}_{2,0}$ is spanned by the top class $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$, and $\mathcal{F}_{2,2}={\mathcal{Q}}_{m}(2)^{G}$. Both are evidently $\mathcal{A}$-submodules of ${\mathcal{Q}}_{m}(2)^{G}$.
Proposition 8.5.
$\mathcal{F}_{3,1}$ is a submodule of ${\mathcal{Q}}_{m}(3)^{G}$ over the Steenrod algebra.
@@
We need to show that for each $k\geq 1$ and $0\leq a<[m]_{q}$, $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{s})$ is in the span of $\delta_{1}^{3}(1)$ and $\delta_{2}^{2}(\Delta^{m}_{1})$. Note that $\delta_{1}^{3}(1)$ spans the one-dimensional space of ${\mathcal{Q}}_{m}(3)^{G}$ in the top degree $3(q^{m}-1)$.
First of all, if $k\geq q^{m}$, then the degree of $\mathcal{P}^{k}(\delta_{2}^{2}Q_{1,0}^{s})$, which is $q^{m}(q-1)+2(q^{m}-q)+s(q-1)$ is at least $3(q^{m}-1)$ unless $q=2$ and $k=2^{m}$, $s=0$. In this range, ${\mathcal{Q}}_{m}(3)$ is either spanned by $\delta_{3}^{3}(1)$ or zero.
If $q=2$, $k=2^{m}$, $s=0$, then $\mathcal{P}^{2^{m}}(\delta_{2}^{2}(1))$ is in degree $13(2^{m}-1)-$ 1. We claim that there is no $G$-invariant polynomial in this degree. Indeed, the only non-zero symmetric polynomial in this degree is
$x_{1}^{2^{m}-1}x_{2}^{2^{m}-1}x_{3}^{2^{m}-2}+x_{1}^{2^{m}-1}x_{2}^{2^{m}-2}x_{3}^{2^{m}-1}+x_{1}^{2^{m}-2}x_{2}^{2^{m}-1}x_{3}^{2^{m}-2}.$
By direct inspection, this polynomial is not invariant under the action of the operation mapping $x_{1}\mapsto x_{1}+x_{2}$ and fixes $x_{2},x_{3}$.
Now let us consider the case $k<q^{m}$. By Lemma 8.3, we have the following equality in $S$:
(8.2) $\mathcal{P}^{k}(\delta_{2}f)+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}f)+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}f)=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}f)+\delta_{2}(\mathcal{P}^{k}f)$
for all polynomial $f$ which is $GL_{1}$-invariant in the first variable. Applying this for $f=Q_{1,0}^{s}$, it follows by induction on $k$ that for $0\leq k<q^{m}$, $\mathcal{P}^{k}(\delta_{2}Q_{1,0}^{s})$ is a linear combination of polynomials of the form
$Q_{2,1}^{a}Q_{2,0}^{b}\delta_{2}(Q_{1,0}^{c}).$
Now applying (8.2) again but for $f=\delta_{2}(Q_{1,0}^{s})$, we get
(8.3) $\mathcal{P}^{k}(\delta_{2}^{2}(Q_{1,0}^{s}))+Q_{2,1}\mathcal{P}^{k-q}(\delta_{2}^{2}(Q_{1,0}^{s}))+Q_{2,0}\mathcal{P}^{k-q-1}(\delta_{2}^{2}(Q_{1,0}^{s}))=\delta_{2}(Q_{1,0}\mathcal{P}^{k-1}\delta_{2}(Q_{1,0}^{s}))+\delta_{2}(\mathcal{P}^{k}\delta_{2}(Q_{1,0}^{s})).$
The right-hand side is then a sum of polynomials of the following forms
  • $\delta_{2}(Q_{1,0}Q_{2,1}^{a}Q_{2,0}^{b}\delta_{2}(Q_{1,0}^{c}))$, or
  • $\delta_{2}(Q_{2,1}^{a}Q_{2,0}^{b}\delta_{2}(Q_{1,0}^{c}))$.
Up to this point, we have been working in $S$. Now we consider the situation in ${\mathcal{Q}}_{m}(3)$. From Lemma 7.2, we see that this sum belongs to the subspace of ${\mathcal{Q}}_{m}(3)$ spanned by polynomials of the forms
  • $\delta_{2}^{2}(Q_{1,0}^{a})$, or
  • $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{a(q-1)}$.
Both sides of Equation 8.3 now are $P(2,1)$-invariant. Take the transfer from $P(2,1)$ to $G$. We have $\mathcal{P}^{i}(\delta_{2}^{2}(Q_{1,0}^{s}))$ is $G$-invariant in ${\mathcal{Q}}_{m}(3)$ and
$\operatorname{Tr}_{P(2,1)}^{G}(Q_{2,1})=\operatorname{Tr}_{P(2,1)}^{G}(Q_{2,0})=\operatorname{Tr}_{P(2,1)}^{G}(x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{a(q-1))})=0.$
The first two terms vanish in $S$ for degree reasons: there are no genuine $G$-invariants in degrees $q^{2}-q$ and $q^{2}-1$. The third vanishes in ${\mathcal{Q}}_{m}(3)$ because if $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}x_{3}^{a(q-1)}$ is a nontrivial summand of a $G$-invariant polynomial in ${\mathcal{Q}}_{m}(3)$, then it is a $P(1,2)$-invariant as well, but from Remark 6.1, this can only happen at the top degree.
It follows that $\mathcal{P}^{k}(\delta_{2}^{2}(Q_{1,0}^{s}))$ must be a scalar multiple of some $\delta_{2}^{2}(Q_{1,0}^{a})$, and thus belongs to $\mathcal{F}_{3,1}$. ∎
@@
If $f$ is a polynomial which is $G_{2}$-invariant in the first two variables, then for any $k\geq 0$, the following holds in $S$:
$\mathcal{P}^{k}(\delta_{3}f)+Q_{3,2}\mathcal{P}^{k-q^{2}}(\delta_{3}f)+Q_{3,1}\mathcal{P}^{k-q^{2}-q}(\delta_{3}f)+Q_{3,0}\mathcal{P}^{k-q^{2}-q-1}(\delta_{3}f)=\\ \delta_{3}\bigg{(}\mathcal{P}^{k}(f)+Q_{2,1}\mathcal{P}^{k-q}(f)+Q_{2,0}\mathcal{P}^{k-q-1}(f)\bigg{)}+\delta_{3;m+1}\bigg{(}\mathcal{P}^{k-q^{m}}(f)+Q_{2,1}\mathcal{P}^{k-q^{m}-q}(f)+Q_{2,0}\mathcal{P}^{k-q^{m}-q-1}(f)\bigg{)}.$
@@
The above equality is again an application of the Cartan formula to $\mathcal{P}^{k}(L_{3}\delta_{3}f)$, using the formulae for the action of the Steenrod powers on $L_{2}$ given in the proof of Lemma 8.3, and similar formulae for these actions on $L_{3}$: $\mathcal{P}^{q^{2}}L_{3}=Q_{3,2}L_{3}$, $\mathcal{P}^{q^{2}+q}L_{3}=Q_{3,1}L_{3}$, $\mathcal{P}^{q^{2}+q+1}L_{3}=Q_{3,0}L_{3}$ and $\mathcal{P}^{k}L_{3}=0$ for all other $k\geq 1$. ∎
The next lemma discusses how to deal with the last term in the above formula.
@@
Recall from the fundamental equation that $X^{q^{3}}-Q_{3,2}X^{q^{2}}+Q_{3,1}X^{q}-Q_{3,0}X=0$ if $X$ is in the span of $x_{1},x_{2},x_{3}$. It follows that
$x_{i}^{q^{m+1}}-Q_{3,2}^{q^{m-2}}x_{i}^{q^{m}}+Q_{3,1}^{q^{m-2}}x_{i}^{q^{m-1}}-Q_{3,0}^{q^{m-2}}x_{i}^{q^{m-2}}=0,\quad 1\leq i\leq 3,$
and we obtain the following identity in $S$ relating $\delta_{3}$ for various $m$:
@@
Note that the degree of $\delta_{3;m+1}(h)$ equals $q^{m+1}-q^{2}+\deg(h)$. If $q>3$, this is strictly greater than $3(q^{m}-1)$, and we can conclude immediately that $\delta_{3;m+1}(h)=0$ in $\mathcal{Q}_{m}(3)$.
Consider now the case $q=3$. Observe that, by direct computation, $Q_{3,0}=Q_{3,1}=0$, $Q_{3,2}=(x_{1}x_{2}x_{3})^{3^{2}-3^{1}}$ in ${\mathcal{Q}}_{2}(3)$. Hence $Q_{3,1}^{q^{m-2}}=Q_{3,0}^{q^{m-2}}=0$ in ${\mathcal{Q}}_{m}(3)$ and the above equation becomes
(8.4) $\delta_{3;m+1}(h)=Q_{3,2}^{q^{m-2}}\delta_{3;m}(h)=(x_{1}x_{2}x_{3})^{3^{m}-3^{m-1}}\delta_{3;m}(h)\quad\text{in ${\mathcal{Q}}_{m}(3)$}.$
On the other hand, for degree reasons, we must have $\deg(h)\leq 3(3^{m}-1)-(3^{m+1}-3^{2})=6$. In this range, there are (up to a scalar) only two Dickson polynomials, namely $h=1$ or $h=Q_{2,1}$. If $h=1$, then from Proposition 2.7, we have immediately
$\delta_{3;m+1}(1)=Q_{3,2}^{3^{m-2}}\delta_{3;m}(1)=\delta_{3;m}(Q_{2,1}^{3^{m-1}}).$
If $h=Q_{2,1}$, we prove by induction on $m\geq 2$ that
$\delta_{3;m+1}(Q_{2,1})=0\quad\text{in ${\mathcal{Q}}_{m}(3)$}.$
Indeed, the base case can be checked directly since $\delta_{3;3}(Q_{2,1})=Q_{3,1}$ in $S$, and the induction step follows immediately from Equation (8.4).
The next and final case is $q=2$. Using the identity
$x_{1}^{2^{m+1}}=x_{1}^{2^{m}}(x_{1}^{2^{m}}+x_{2}^{2^{m}}+x_{3}^{2^{m}})-x_{1}^{2^{m}}(x_{2}^{2^{m}}+x_{3}^{2^{m}}),$
and similarly for $x_{2}^{2^{m+1}}$ and $x_{3}^{2^{m+1}}$, we see that $\delta_{3;m+1}(h)$ equals in ${\mathcal{Q}}_{m}(3)$ to
@@
It is straightforward to verify the following equality in $S$:
@@
We then obtain an equality in ${\mathcal{Q}}_{m}(3)$:
$\delta_{3;m+1}(h)=\delta_{3;m}(Q_{2,1}h\delta_{2;m}(1))-\delta_{3;m}(h\delta_{2;m}(Q_{1,0}^{2})).$
The conclusion follows from Corollary 7.3. ∎
@@
With a little more work, using 7.2, it can be shown that when $q=2$, then for $h\in{\mathcal{D}}_{2}$, we have the following explicit formula for $\delta_{3;m+1}(h)$ in ${\mathcal{Q}}_{m}(3)$:
$\delta_{3;m+1}(h)=\begin{cases}0&\text{if $h$ is a multiple of $Q_{2,0}$;}\\ \delta_{2;m}^{2}(Q_{1,0}^{2s})&\text{if $h=Q_{2,1}^{s}$, $s\geq 0$}.\end{cases}$
@@
Now we can finish the proof of Theorem 1.12. It remains to show that $\mathcal{F}_{3,2}$ is an $\mathcal{A}$-submodule of ${\mathcal{Q}}_{m}(3)^{G}$. From Lemma 8.7, we see that if $f\in{\mathcal{D}}_{2}$, then
$\mathcal{P}^{k}(\delta_{3}f)+Q_{3,2}\mathcal{P}^{k-q^{2}}(\delta_{3}f)+Q_{3,1}\mathcal{P}^{k-q^{2}-q}(\delta_{3}f)+Q_{3,0}\mathcal{P}^{k-q^{2}-q-1}(\delta_{3}f)\in\mathcal{F}_{3,2}.$
We can proceed by induction. If $\mathcal{P}^{i}(\delta_{3}f)\in\mathcal{F}_{3,2}$ for all $i<k$, then $Q_{3,j}\mathcal{P}^{i}(\delta_{3}f)$ also belongs to $\mathcal{F}_{3,2}$ by Proposition 2.7. It follows that $\mathcal{P}^{k}(\delta_{3}f)$ also belongs to $\mathcal{F}_{3,2}$ and the proof is complete.